# Training Verification-Friendly Neural Networks via Neuron Behavior Consistency

### v1

@@ Abstract
Formal verification provides critical security assurances for neural networks, yet its practical application suffers from the long verification time. This work introduces a novel method for training verification-friendly neural networks, which are robust, easy to verify, and relatively accurate. Our method integrates neuron behavior consistency into the training process, making neuron activation states consistent across different inputs in a local neighborhood, reducing the number of unstable neurons and tightening the bounds of neurons thereby enhancing neural network verifiability. We evaluated our method using the MNIST, Fashion-MNIST, and CIFAR-10 datasets across various network architectures. The results of the experiment demonstrate that networks trained using our method are verification-friendly across different radii and different model architectures, whereas other tools fail to maintain verifiability as the radius increases. We also show that our method can be combined with existing methods to further improve the verifiability of networks.
@@ Introduction
Neural networks are increasingly applied in safety-critical domains such as autonomous driving (Urmson and Whittaker 2008) and flight control (Julian, Kochenderfer, and Owen 2019). However, they frequently struggle with a lack of robustness, as even minor perturbations to their inputs can lead to incorrect predictions (Bu et al. 2022; Chen et al. 2021; Song et al. 2021; Chen et al. 2023, 2022b; Zhao et al. 2021; Chen et al. 2022a). This is unacceptable in safety-critical applications, where consistent performance is imperative. Thus, it is desirable to develop methods to systematically advance the robustness verification of neural networks.
@@
Existing methods for analyzing robustness can be divided into two categories: empirical analysis through adversarial attacks and mathematical proof via formal verification. While adversarial attacks generate misleading examples, they merely demonstrate the presence of adversarial samples without affirming their absence. In contrast, formal verification ensures the correctness of neural networks using logical and mathematical methods, allowing us to formally verify the robustness of neural networks. This rigorous verification is essential for safety-critical systems.
@@
Advanced formal verification tools (Wang et al. 2021; Zhang et al. 2022a; Bak 2021; Katz et al. 2017), typically employ branch-and-bound algorithms for neural network verification. At the beginning of verification, abstract interpretation (Gehr et al. 2018; Mirman, Gehr, and Vechev 2018; Zhang et al. 2021, 2023; Guo et al. 2021) is usually used to abstract neurons. If the properties of the network remain undetermined after using symbolic propagation (Singh et al. 2019) to calculate neuron boundaries and using MILP (Tjeng, Xiao, and Tedrake 2019; Tran et al. 2020; Zhang et al. 2022b; Zhang, Song, and Sun 2023; Zhang et al. 2024) or SMT methods (Ehlers 2017; Huang et al. 2017; Katz et al. 2017; Zhao et al. 2022; Liu et al. 2024) for constraint solving, further branching is required. The branching process involves enumerating the activation states of unstable neurons, whose activation status cannot be determined through bound calculations, to introduce additional constraints, thereby refining the abstraction. Typically, neural networks contain numerous unstable neurons, and exploring these combinations of activation states requires exponential time, which limits the widespread application of formal verification techniques in practice.
@@
In addition to developing ever-more sophisticated methods for post-training verification, researchers have investigated the idea of training neural networks that are easier to verify, as known as verification-friendly neural networks. Ideally, a training method for verification-friendly neural networks must satisfy the following requirements. First, (accuracy) the resultant neural network must have an accuracy that is comparable to neural networks trained conventionally. Second, (robustness) the resultant neural network must have improved robustness, which could be measured using existing adversarial attacks. Third, (verifiability) it must be easier to verify using existing or dedicated neural network verification techniques, which can be measured using the effectiveness of selected neural network verification methods.
@@
There are mainly two existing approaches to promoting the verification-friendliness of neural networks. One approach involves post-processing the network through methods that modify the weights of the networks (Xiao et al. 2019; Baninajjar, Rezine, and Aminifar 2023). Another involves altering the neural network design and training process with considerations for verification (Xiao et al. 2019; Narodytska et al. 2019; Xu et al. 2024). However, these methods still have limitations. The effectiveness of post-training methods is limited by the network itself. Moreover, certain post-training methods such as using MILP to make the network more sparse (Baninajjar, Rezine, and Aminifar 2023), are computationally expensive and may not be adaptable for large networks. The ReLU Stable methods, which introduce ReLU Stable (RS) loss (Xiao et al. 2019) to reduce the number of unstable neurons, depend on the bounds of neurons, and when the perturbation radius changes, it often causes these bounds to shift, affecting the verification efficiency of the network. Certified training, mainly based on the heavy Interval Bound Propagation (IBP) method (Mirman, Gehr, and Vechev 2018; Gowal et al. 2018; Zhang et al. 2019a; Xu et al. 2020) suffers from a long training time and the gradient explosion or vanish problem. Adversarial training (Madry et al. 2018; Zhang et al. 2019b; Ganin et al. 2016; Zhu et al. 2017) typically increase network complexity but do not contribute to improving network verifiability.
@@
In this work, we introduce a straightforward yet effective training method that enhances the verifiability of neural networks by enforcing the consistency of neuronal behavior, which we refer to as neuron behavior consistency (NBC), throughout the training process as a regularization term. A neuron is called behavior consistent if its activation state remains the same in a given input neighborhood. By maximizing the consistency of neurons, the unstable neurons are decreased, reducing the search space of the verification process. NBC can also help to tight the the bounds of neurons, as fewer unstable neurons introduce less error during bound calculation algorithms. Our approach can be scaled to larger networks compared to MILP-based methods. Moreover, the core of our method lies in the consistency of neuronal behavior without relying on heavy IBP methods, making a reduction in training epochs and ensuring the trained network maintains verifiability across different perturbation radii.
@@
We evaluate our method using Fashion-MNIST, MNIST, and CIFAR-10 datasets across different architectures at various perturbation radii. Our method oputperforms in stable neuron ratio and achieved up to a speedup of 450% in verification time. Importantly, our method accelerates the verification while preserving the accuracy of the model , which is not commonly achieved by existing methods. In summary, our contributions are as follows:
@@
  • We evaluated our methodology on three well-known datasets. Experiments show that networks trained using our method can maintain verification-friendly properties across different radii and different model architectures.
@@
where $\mathcal{C}\subseteq\mathbb{R}^{m_{\mathrm{in}}}$ is the input constraints and $\mathcal{P}\subseteq\mathbb{R}^{m_{\mathrm{out}}}$ is the output constraints.
We focus on the local robustness verification problem $\langle f,\mathcal{C}_{\epsilon}(\boldsymbol{x}),\mathcal{P}_{c}\rangle$, which checks if the classification result $c$ is robust to input perturbations $\varepsilon$ in the norm $l_{\infty}$, where the input constraints $\mathcal{C}_{\epsilon}(\boldsymbol{x})$ defined as $\{\boldsymbol{x}\in\mathbb{R}^{m_{\mathrm{in}}}\mid|\boldsymbol{x}-\boldsymbol{x}^{0}|_{\infty}\leq\varepsilon\}$ and the output constraints $\mathcal{P}_{c}$ defined as $\{\boldsymbol{y}\in\mathbb{R}^{m_{\mathrm{out}}}\mid\bigwedge_{i\neq c}\boldsymbol{y}_{i}-\boldsymbol{y}_{c}\leq 0\}$.
@@
As shown in Figure 1, the verification process starts with the bound calculation, where the upper and lower bounds of neuron output are estimated under specified input constraints. If these bounds are sufficiently precise, the properties of the network can be directly verified. Due to non-linear activation functions such as ReLU, computing these bounds can be complex, which is a problem often addressed through neuron-wise abstraction (Singh et al. 2019; Bak 2021). This abstraction uses linear bounds to approximate neuron outputs, thereby simplifying the bound calculations.
When the bound calculation does not suffice to verify the network’s properties, constraint solving is employed. This process involves determining if the constraints of the abstracted network can be satisfied, using Linear Programming (LP) or Mixed Integer Linear Programming (MILP) methods. Constraints typically include the input constraints $\mathcal{C}$, the negated output constraints $\neg\mathcal{P}$ and the constraints of the abstracted network itself. If these constraints are UNSAT (unsatisfiable), the property holds, which proves that the network can not be successfully attacked within $\mathcal{C}$; otherwise, the returned counterexample must be examined. If the counterexample is not a false positive caused by over-approximation, a violation of the network’s properties is identified. Otherwise, it suggests that the network is over-abstracted, necessitating branch selection to refine the abstraction.
Branch selection selects an unstable neuron (whose activation state cannot be determined through bound calculations) and splits it into active and inactive branches. Identifying the state of a neuron entails establishing new constraints that refine the abstraction. Branch selection strategies critically affect verification efficiency. While branching may require exploration of all unstable neurons in worst-case scenarios, minimizing these neurons reduces the theoretical branching upper bound.
@@
The ordinary training objective does not impose constraints on neurons, potentially resulting in a large number of unstable neurons. We thus propose an alternative training objective that aims to maximize the consistency of neurons. A neuron is consistent if its activation state remains the same in a given input neighborhood. Formally, given an input $\boldsymbol{x}$ and a neighboring input $\boldsymbol{x}^{\prime}$, the consistency of the $j$-th neuron of the $i$-th layer $n^{(i)}_{j}$ is defined as:
@@
The consistency metric presented in Equation 3 is a discrete measure that cannot be directly integrated into the loss function. Therefore, we employ a continuous metric to approximate the consistency of neural behavior across different inputs within the neighborhood domain, which is shown in Algorithm 1. This algorithm calculates the NBC between the neural network $f$ when fed with $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$.
@@
During the calculation of NBC, the NBC value $s$ is initially set to zero, then iterating over the hidden layers of $f$, the algorithm assesses the consistency for each neuron between the original and adversarial images, incrementally updating NBC. The final NBC is calculated as the sum of the scaled neuron consistencies across all layers.
Due to the characteristics of gradient backpropagation, the layer close to the output layer has a greater impact on the network’s behavior. Therefore, we use the KL divergence as a consistency metric for the output layer as a regularization term to ensure that the network’s output remains consistent across different inputs, which can be calculated as:
@@
For the hidden layers, to prevent gradient explosion or vanishing, we use cosine similarity as a continuous metric to approximate the consistency of behavior, in which the output is in the range of [0,1], making it more suitable for training neural networks with high-dimensional intermediate layers. To balance the impact of layers with different numbers of neurons, we scale the NBC value by driving $\gamma[i]$ at line 6 of Algorithm 1. Our intuition is to prioritize layers with smaller dimensionsas their consistent behavior may may constrain subsequent layers through propagation. Moreover, input and output layers often have fewer neurons, applying constraints to them can directly affect the forward and backward propagation process and accelerate the convergence of target loss. In contrast, over-constraining middle layers, which have more neurons and extract complex features, could limit the model’s expressive power. Roughly speaking, using a looser penalty on layers with more neurons may slightly decrease accuracy but increase the proportion of stable neurons and improve verification speed. In our experiments, we apply factor $\gamma[i]=2^{r[i]}$ to balance the impact of layers with different numbers of neurons, where $r[i]$ means the number of neurons in the $i$-th hidden layer is the $r[i]$-th smallest number in all the numbers of neurons in hidden layers.
@@
1:  Generate a random starting point $\boldsymbol{x}^{\prime}\in\mathcal{S}(\boldsymbol{x},\varepsilon)$
@@
Algorithm 2 outlines the NBC loss calculation process. This algorithm uses the idea of adversarial training to find an adversarial input $\boldsymbol{x}^{\prime}$ that minimizes the NBC loss between the input $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$. The algorithm starts by randomly selecting an adversarial input $\boldsymbol{x}^{\prime}$ within the $\varepsilon$-neighborhood of the original input $\boldsymbol{x}$. Subsequently, the adversarial input undergoes iterative perturbation over $k$ steps, with each perturbation step adjusting the adversarial input by gradients of the NBC for the adversarial input, scaled by the step size $\alpha$. The adversarial input is clipped after each perturbation step to ensure that it remains within the perturbation range allowed from the original input. The final NBC loss is calculated as the summation of the cross-entropy loss and the negation of scaled NBC (Algorithm 1) between the original input and the adversarial input. Detailed disgussion of the hyperparameters $\gamma$ in Algorithm 1 and $\beta$ in Algorithm 2 are provided in supplementary material.
We train the network to maximize neural response uniformity across varied inputs within the neighborhood, thereby reducing the number of unstable neurons and tightening neuronal bounds. Consequently, the network trained with NBC loss is more friendly to formal verification methods, as it exhibits fewer unstable neurons and more precise bounds.
@@
In this section, we evaluate our methods to answer the following research questions:
RQ1: Can networks trained with our method maintain their verification-friendly properties across various network architectures and radii?
@@
RQ3: How does our method compare with existing methods under close accuracy ranges?
@@
Dataset. Networks are trained on three widely used datasets, MNIST (LeCun et al. 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR-10 (Krizhevsky 2009).
Network Architecture. We select neural networks of varying sizes from VNN-COMP (Bak, Liu, and Johnson 2021; Müller et al. 2023) to assess the effectiveness of the methods used. For the MNIST and Fashion-MNIST datasets, we chose the M1 (cnn_4_layer), M2 (relu_stable), and M3 (conv_big) models, with approximately 0.16M, 0.17M, and 1.9M parameters, respectively. For the CIFAR-10 dataset, we select the C1 (marabou_medium), C2 (marabou_large), and C3 (conv_big) models, containing about 0.17M, 0.34M, and 2.4M parameters, respectively. Network architectures are detailed in supplementary material.
Baselines. We choose Relu Stable (Xiao et al. 2019) as a baseline, which shares the most similarities with our approach. TRADES (Zhang et al. 2019b) and Madry (Madry et al. 2018), two commonly used adversarial training methods, are selected to show that directly using our method can at least achieve robustness comparable to classical robust training methods and combining our method with existing methods can improve verification performance.
Training. We set the batch size to 128, the learning rate to 0.001, and employing the Adam optimizer. For RQ1, networks are trained under default settings for 400 epochs. For RQ2, each network underwent 200 epochs of training with the original method followed by another 200 epochs using a combination of the original and our methods, or in the reverse order. For RQ3, we use CE loss to train a base model and further fine-tune using RS, Madry, Trades, and our method separately to maintain each method’s accuracy within a specified range.
Verification. We use $\alpha,\beta$-CROWN , a state-of-the-art verification tool that performs overall best in VNN-COMP competitions, to verify the properties of the trained networks. For each dataset, we select $k$ images from each of the 10 categories in the test set. For each image $\boldsymbol{x}$, and ground truth label $y$, we verified the property that the network’s output label remain $y$ for input $x$ under each perturbation $\varepsilon$. We set $k=100$ and timeout at 120 seconds for MNIST and Fashion-MNIST, and $k=20$ and timeout at 180 seconds for CIFAR-10.
@@
  • UNSAT%: The percentage of properties that are verified as holds (UNSAT), providing insight into the network’s overall verification effectiveness and verified robustness.
@@
We observe that our method outperforms others in terms of verification time, stable neuron ratio, and verified ratio (UNSAT%). While our method slightly lags in accuracy and PGD accuracy, it generally maintains comparable accuracy to other methods. This result is a comprehensive evaluation across various perturbation radii and models, indicating that networks of various architectures trained using our method are more verification-friendly at different perturbation radii than those trained using other methods.
Table 1 shows the detailed results of networks trained on MNIST. Detailed results of Fashion-MNIST and CIFAR-10 are provided in the supplementary material. As shown in Table 1, traditional adversarial training methods enhance network robustness, yet verifiability declines as the perturbation radius increases. This indicates the need for new training techniques that support effective verification. Regarding UNSAT%, our method consistently outperforms others, especially at higher perturbation radii. For instance, when $\varepsilon=0.2$, our verified ratio reached 46.3%, which is more than 20 times that of the next best-performing method and over 231 times that of the least effective method on the M3 model. Unlike other methods, where UNSATsignificantly drops as the radius increases, our method maintains a high verified ratio across all tested radii.
As for stable neurons, our method excels across nearly all models and perturbation settings. Notably, in the M3 model at $\varepsilon=0.3$, while competing methods exhibited virtually no stable neurons, our method preserved a stable neuron ratio above 50%. A higher proportion of stable neurons means if the verification time is increased, our method is more likely to be verified (as the search space is significantly reduced).
Regarding average verification time, including UNSAT and Timeout issues, our method generally requires less time across all verification tasks. An exception occurs in the M3 model at a 0.3 perturbation radius, where verification failed to confirm all the properties of our model. Conversely, the model trained using the Madry has many properties quickly verified as violated, thus shortening the times. Furthermore, under a 100-step PGD attack, our method achieved comparable accuracy to these adversarial training approaches. Such results also hold for Fashion-MNIST and CIFAR-10 datasets (see supplementary material).
@@
Answer RQ1: Our method maintains high stable neuron ratios, UNSAT%, and robustness across various perturbation radii. Overall, networks trained with our method maintain their verification-friendly properties across various network architectures and radii.

### v2

@@ Abstract
Formal verification provides critical security assurances for neural networks, yet its practical application suffers from the long verification time. This work introduces a novel method for training verification-friendly neural networks, which are robust, easy to verify, and relatively accurate. Our method integrates neuron behavior consistency into the training process, making neuron activation states remain consistent across different inputs within a local neighborhood. This reduces the number of unstable neurons and tightens the bounds of neurons thereby enhancing the network’s verifiability. We evaluated our method using the MNIST, Fashion-MNIST, and CIFAR-10 datasets with various network architectures. The experimental results demonstrate that networks trained using our method are verification-friendly across different radii and architectures, whereas other tools fail to maintain verifiability as the radius increases. Additionally, we show that our method can be combined with existing approaches to further improve the verifiability of networks.
@@ Introduction
Neural networks are increasingly being applied in safety-critical domains such as autonomous driving (Urmson and Whittaker 2008) and flight control (Julian, Kochenderfer, and Owen 2019). However, they often struggle with a lack of robustness, as even minor perturbations to their inputs can lead to incorrect predictions (Bu et al. 2022; Chen et al. 2021; Song et al. 2021; Chen et al. 2023, 2022b; Zhao et al. 2021; Chen et al. 2022a). This is unacceptable in safety-critical applications, where consistent performance is imperative. Thus, it is desirable to develop methods to systematically advance the robustness verification of neural networks.
@@
Existing methods for analyzing robustness can be broadly classified into two categories: empirical analysis through adversarial attacks and mathematical proof via formal verification. While adversarial attacks generate misleading examples, they merely demonstrate the presence of adversarial samples without affirming their absence. In contrast, formal verification ensures the correctness of neural networks using logical and mathematical methods, allowing us to verify their robustness formally. This rigorous verification is essential for safety-critical systems.
@@
Advanced formal verification tools (Wang et al. 2021; Zhang et al. 2022a; Bak 2021; Katz et al. 2017), typically employ branch-and-bound algorithms for neural network verification. At the start of the verification process, abstract interpretation (Gehr et al. 2018; Mirman, Gehr, and Vechev 2018; Zhang et al. 2021, 2023; Guo et al. 2021) is usually used to abstract neurons. If the properties of the network remain undetermined after using symbolic propagation (Singh et al. 2019) to calculate neuron boundaries and applying MILP (Tjeng, Xiao, and Tedrake 2019; Tran et al. 2020; Zhang et al. 2022b; Zhang, Song, and Sun 2023; Zhang et al. 2024) or SMT methods (Ehlers 2017; Huang et al. 2017; Katz et al. 2017; Zhao et al. 2022; Liu et al. 2024a) for constraint solving, further branching is required. The branching process involves enumerating the activation states of unstable neurons, whose activation status cannot be determined through bound calculations. This introduces additional constraints, thereby refining the abstraction. Typically, neural networks contain numerous unstable neurons, and exploring the combinations of these activation states requires exponential time, which limits the widespread application of formal verification techniques in practice.
@@
In addition to developing ever-more sophisticated methods for post-training verification, researchers have investigated the idea of training neural networks that are easier to verify, known as verification-friendly neural networks. Ideally, a training method for verification-friendly neural networks must satisfy the following requirements. First, (accuracy) the resultant neural network must have an accuracy comparable to that of neural networks trained conventionally. Second, (robustness) the resultant neural network must have improved robustness, which could be measured using existing adversarial attacks. Third, (verifiability) it must be easier to verify using existing or dedicated neural network verification techniques, which can be measured using the effectiveness of selected neural network verification methods.
@@
There are mainly two existing approaches to promoting the verification-friendliness of neural networks. One approach involves post-processing the network using methods that modify the weights of the networks (Xiao et al. 2019; Baninajjar, Rezine, and Aminifar 2024). The other involves altering the neural network design and training process with considerations for verification (Xiao et al. 2019; Narodytska et al. 2019; Xu et al. 2024). However, these methods still have limitations. The effectiveness of post-training methods is limited by the network itself. Additionally, certain post-training methods such as using MILP to make the network sparser (Baninajjar, Rezine, and Aminifar 2024), are computationally expensive and may not scale well for large networks. The ReLU Stable methods, which introduce ReLU Stable (RS) loss (Xiao et al. 2019) to reduce the number of unstable neurons, depend on the bounds of neurons. When the perturbation radius changes, it often causes these bounds to shift, affecting the verification efficiency of the network. Certified training, which is primarily based on the heavy Interval Bound Propagation (IBP) method (Mirman, Gehr, and Vechev 2018; Gowal et al. 2018; Zhang et al. 2019a; Xu et al. 2020) suffers from long training times and issues like gradient explosion or vanish problem. Adversarial training (Madry et al. 2018; Zhang et al. 2019b; Ganin et al. 2016; Zhu et al. 2017) typically increases network robustness but does not contribute to improving its verifiability.
@@
In this work, we introduce a straightforward yet effective training method that enhances the verifiability of neural networks by enforcing the consistency of neuron behavior, which we refer to as neuron behavior consistency (NBC), throughout the training process as a regularization term. A neuron is called behavior consistent if its activation state remains the same within a given input neighborhood. By maximizing the consistency of neurons, the unstable neurons are decreased, reducing the search space of the verification process. NBC can also help tighten the bounds of neurons, as fewer unstable neurons introduce less error during bound calculation algorithms. Our approach can be scaled to larger networks compared to MILP-based methods. Moreover, the core of our method lies in the consistency of neuron behavior without relying on heavy IBP methods, which reduces training epochs and ensures that the trained network maintains verifiability across different perturbation radii.
@@
We evaluate our method using Fashion-MNIST, MNIST, and CIFAR-10 datasets across different architectures at various perturbation radii. Our method outperforms others in stable neuron ratio and achieves up to a 450% speedup in verification time. Importantly, our method accelerates the verification while preserving the accuracy of the models, which is not commonly achieved by existing methods. In summary, our contributions are as follows:
@@
  • We evaluate our method on three well-known datasets. Experimental results show that networks trained using our method can maintain verification-friendly properties across different radii and different model architectures.
@@
where $\mathcal{C}\subseteq\mathbb{R}^{m_{\mathrm{in}}}$ represents the input constraints and $\mathcal{P}\subseteq\mathbb{R}^{m_{\mathrm{out}}}$ represents the output constraints.
We focus on the local robustness verification problem $\langle f,\mathcal{C}_{\epsilon}(\boldsymbol{x}),\mathcal{P}_{c}\rangle$, which checks if the classification result $c$ is robust to input perturbations $\varepsilon$ in the $l_{\infty}$-norm, where the input constraints $\mathcal{C}_{\epsilon}(\boldsymbol{x})$ are defined as $\{\boldsymbol{x}\in\mathbb{R}^{m_{\mathrm{in}}}\mid|\boldsymbol{x}-\boldsymbol{x}^{0}|_{\infty}\leq\varepsilon\}$ and the output constraints $\mathcal{P}_{c}$ are defined as $\{\boldsymbol{y}\in\mathbb{R}^{m_{\mathrm{out}}}\mid\bigwedge_{i\neq c}\boldsymbol{y}_{i}-\boldsymbol{y}_{c}\leq 0\}$.
@@
As shown in Figure 1, the verification process begins with the bound calculation, where the upper and lower bounds of neuron outputs are estimated under specified input constraints. If these bounds are sufficiently precise, the properties of the network can be directly verified. Due to non-linear activation functions such as ReLU, computing these bounds can be complex, which is a problem often addressed through neuron-wise abstraction (Singh et al. 2019; Bak 2021). This abstraction uses linear bounds to approximate neuron outputs, thereby simplifying the bound calculations.
When the bound calculation does not suffice to verify the network’s properties, constraint solving is employed. This process involves determining if the constraints of the abstracted network can be satisfied, using Linear Programming (LP) or Mixed Integer Linear Programming (MILP) methods. Constraints typically include the input constraints $\mathcal{C}$, the negated output constraints $\neg\mathcal{P}$, and the constraints of the abstracted network itself. If these constraints are UNSAT (unsatisfiable), the property holds, proving that the network can not be successfully attacked within $\mathcal{C}$; otherwise, the returned counterexample must be examined. If the counterexample is not a false positive caused by over-approximation, it indicates a violation of the network’s properties. Otherwise, it suggests that the network is over-abstracted, necessitating branch selection to refine the abstraction.
Branch selection selects an unstable neuron (whose activation state cannot be determined through bound calculations) and splits it into active and inactive branches. Identifying the state of a neuron entails establishing new constraints that refine the abstraction. Branch selection strategies significantly affect verification efficiency. While branching may require the exploration of all unstable neurons in worst-case scenarios, minimizing these neurons can exponentially reduce the theoretical upper bound on branching.
@@
The ordinary training objective does not impose constraints on neurons, potentially resulting in a large number of unstable neurons. We thus propose an alternative training objective that aims to maximize the consistency of neurons. A neuron is called consistent if its activation state remains the same within a given input neighborhood. Formally, given an input $\boldsymbol{x}$ and a neighboring input $\boldsymbol{x}^{\prime}$, the consistency of the $j$-th neuron in the $i$-th layer $n^{(i)}_{j}$ is defined as:
@@
The consistency metric presented in Equation 3 is a discrete measure and cannot be directly integrated into the loss function. Therefore, we employ a continuous metric to approximate the consistency of neural behavior across different inputs within the neighborhood domain, as shown in Algorithm 1. This algorithm calculates the NBC for the neural network $f$ when fed with $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$.
@@
During the calculation of NBC, the NBC value $s$ is initially set to zero. Then iterating over the hidden layers of $f$, the algorithm assesses the consistency of each neuron between the original and adversarial images, incrementally updating the NBC. The final NBC is computed as the sum of the scaled neuron consistencies across all layers.
Due to the characteristics of gradient backpropagation, the layer close to the output layer has a greater impact on the network’s behavior. Therefore, we use the KL divergence as a consistency metric for the output layer, serving as a regularization term to ensure that the network’s output remains consistent across different inputs. This can be calculated as:
@@
For the hidden layers, to prevent gradient explosion or vanishing, we use cosine similarity as a continuous metric to approximate the consistency of behavior. Cosine similarity outputs a value in the range of [0,1], making it more suitable for training neural networks with high-dimensional intermediate layers.
To balance the impact of layers with different numbers of neurons, we scale the NBC value by the factor $\gamma[i]$ at line 6 of Algorithm 1. Our intuition is to prioritize layers with smaller dimensions, as their more consistent behavior can help propagate constraints through the network. Additionally, layers near the input and output often have fewer neurons. Applying constraints to these layers directly influences the forward and backward propagation processes, which can accelerate the convergence of the target loss. In contrast, over-constraining the middle layers, which contain more neurons and extract more complex features, could limit the model’s expressive power. More generally, applying stricter penalties to layers may decrease accuracy but increase the proportion of stable neurons. For this reason, we mainly impose constraints on layers with fewer neurons. In our experiments, we apply the factor $\gamma[i]=2^{r[i]}$ to balance the impact of layers with different numbers of neurons, where $r[i]$ means the number of neurons in the $i$-th hidden layer $m[i]$ is the $r[i]$-th smallest number in all the numbers of neurons in hidden layers.
@@
1:  Generate a random starting point $\boldsymbol{x}^{\prime}\in\mathcal{C}_{\epsilon}(\boldsymbol{x})$
@@
Algorithm 2 outlines the NBC loss calculation process. This algorithm uses the idea of adversarial training to find an adversarial input $\boldsymbol{x}^{\prime}$ that minimizes the NBC loss between the input $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$. The algorithm starts by selecting a random adversarial input $\boldsymbol{x}^{\prime}$ within the $\varepsilon$-neighborhood of the original input $\boldsymbol{x}$. The adversarial input is then iteratively perturbed over $k$ steps, with each perturbation adjusting the adversarial input based on the gradients of the NBC concerning $\boldsymbol{x}^{\prime}$ , scaled by the step size $\alpha$. The adversarial input is clipped after each perturbation step to ensure it remains within the perturbation range allowed by the original input. The final NBC loss is computed as the summation of the cross-entropy loss and the negation of scaled NBC (calculated according to Algorithm 1) between the original input and the adversarial input. A detailed discussion of the hyperparameters $\gamma$ in Algorithm 1 and $\beta$ in Algorithm 2 is provided in supplementary material (Liu et al. 2024b).
We train the network to maximize the consistency of neuron behavior across varied inputs within the neighborhood, thereby reducing the number of unstable neurons and tightening neuron bounds. As a result, the network trained with the NBC loss is friendly to formal verification methods, as it exhibits fewer unstable neurons and more precise bounds.
@@
In this section, we evaluate our method to address the following research questions:
RQ1: Can networks trained with our method maintain verification-friendly properties across various network architectures and perturbation radii?
@@
RQ3: How does the performance of our method compare to existing methods when their accuracies are close?
@@
Dataset. Networks are trained on three widely used datasets: MNIST (LeCun et al. 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR-10 (Krizhevsky 2009).
Network Architecture. We select neural networks of varying sizes from VNN-COMP (Bak, Liu, and Johnson 2021; Müller et al. 2023) to assess the effectiveness of the methods used. For the MNIST and Fashion-MNIST datasets, we chose the M1 (cnn_4_layer), M2 (relu_stable), and M3 (conv_big) models, with approximately 0.16M, 0.17M, and 1.9M parameters, respectively. For the CIFAR-10 dataset, we select the C1 (marabou_medium), C2 (marabou_large), and C3 (conv_big) models, containing about 0.17M, 0.34M, and 2.4M parameters, respectively. Network architectures are detailed in supplementary material (Liu et al. 2024b).
Baselines. We choose Relu Stable (Xiao et al. 2019) as a baseline, as it shares the most similarities with our approach. TRADES (Zhang et al. 2019b) and Madry (Madry et al. 2018), two commonly used adversarial training methods, are selected to show that directly using our method can at least achieve robustness comparable to classical robust training methods and combining our method with existing methods can improve verification performance.
Training. The batch size is set to 128, and using the Adam optimizer. For RQ1, networks are trained under default settings for 400 epochs. For RQ2, each network is trained for 200 epochs using the original method, followed by an additional 200 epochs combining the original method with our approach, or vice versa. For RQ3, we train a base model using the CE loss, then fine-tune it separately using RS, Madry, TRADES, and our method, ensuring that each method maintains accuracy within a specified range.
Verification. We use $\alpha,\beta$-CROWN, a state-of-the-art verification tool that performs the best in VNN-COMP competitions, to verify the properties of the trained networks. For each dataset, we select $k$ images from each of the 10 categories in the test set. For each image $\boldsymbol{x}$ and its ground truth label $y$, we verify the property that the network’s output label remains $y$ for input $\boldsymbol{x}$ under each perturbation $\varepsilon$. We set $k=100$ and a timeout of 120 seconds for MNIST and Fashion-MNIST, and $k=20$ with a timeout of 180 seconds for CIFAR-10.
@@
  • UNSAT%: The percentage of properties verified to hold (UNSAT), indicating the network’s overall verification effectiveness and robustness.
@@
We observe that our method outperforms others in terms of verification time, stable neuron ratio, and verified ratio (UNSAT%). While our method slightly lags in accuracy and PGD accuracy, it generally maintains comparable accuracy to the other methods. This result represents a comprehensive evaluation across various perturbation radii and models, indicating that networks of various architectures trained using our method are more verification-friendly at different perturbation radii than those trained using other methods.
Table 1 shows the detailed results for networks trained on MNIST. Detailed results for Fashion-MNIST and CIFAR-10 are provided in the supplementary material. As shown in Table 1, traditional adversarial training methods enhance network robustness, yet verifiability declines as the perturbation radius increases. This suggests the need for new training techniques that support effective verification.
Regarding UNSAT%, our method consistently outperforms others, especially at higher perturbation radii. For instance, when $\varepsilon=0.2$, our verified ratio reached 46.3%, more than 20 times that of the next best-performing method and over 231 times that of the least effective method on the M3 model. Unlike other methods, where UNSATsignificantly drops as the radius increases, our method maintains a high verified ratio across all tested radii.
As for stable neurons, our method excels across nearly all models and perturbation settings. Notably, in the M3 model at $\varepsilon=0.3$, while competing methods exhibited virtually no stable neurons, our method preserved a stable neuron ratio above 50%. A higher proportion of stable neurons means that if the verification time is increased, our method is more likely to be successfully verified (as the search space is significantly reduced).
Regarding average verification time, including UNSAT and Timeout issues, our method generally requires less time across all verification tasks. An exception occurs in the M3 model at a 0.3 perturbation radius, where verification failed to confirm all properties of our model. In contrast, the model trained using Madry quickly verified many properties as violated, thus shortening the verification times. Furthermore, under a 100-step PGD attack, our method achieved comparable accuracy to these adversarial training approaches. These results also hold for the Fashion-MNIST and CIFAR-10 datasets (see supplementary material).
@@
Answer to RQ1: Our method maintains high stable neuron ratios, UNSAT%, and robustness across various perturbation radii. Overall, networks trained with our method preserve their verification-friendly properties across different network architectures and radii.